OpenStudy (anonymous):

Solutions of x^2+12x+10=0 complete the square

OpenStudy (zepp):

You take the coefficient of the second term, (12x) and multiply it by 1/2, then square the result ((1/2)*12)^2 = 36 So you plug it in (x^2 + 12 x + 36) + 10 - 36 = 0 (x+6)(x+6) + 10 - 36 = 0 (x+6)^2 = 26 Uh, solution is ^ ;D